Status Report

Aqua Status Report May 25, 2002 – May 27, 2002

By SpaceRef Editor
May 27, 2002
Filed under , ,

The following activities were completed:

  1. Uplinked the daily Master Command Loads and Ephemeris
  2. CERES Fore/Aft Az Position & Rate Functional Test
  3. CERES Packet Data Type Functional Test
  4. CERES Bridge Balance Functional Test
  5. CERES XTrack Internal Cal Functional Test
  6. CERES Fore/Aft Scan Timeout Functional Test
  7. AIRS Warm Functional Checkout 1 & 2
  8. AIRS Vis/NIR Radiometric Test Lamp 1,2,3
  9. CERES Elevation Scan Functional Test
  10. CERES Fore/Aft Biaxial Internal Cal Functional Test
  11. CERES Solar Cal Functional Test
  12. CERES Fore/Aft Contamination Safe Functional Test
  13. CERES Hold Sequence Functional Test

The following activities are scheduled for the next 24 hours:

  1. Uplink the daily Master Command Loads and Ephemeris.
  2. CERES XTrack Internal Calibration
  3. CERES Fore/Aft SPS1 Sunrise/Sunset Test
  4. CERES Fore/Aft SPS1/2 Sunrise/Sunset Test 1 & 2
  5. CERES XTrack

Aqua Spacecraft Status:

  1. Spacecraft State is currently Standby.
  2. GNC Mode is Fine Point Mode, HiFi Ephemeris.

Aqua Instrument Status:

  1. AIRS is in Decontaminate state, Quiet bus is ON, Noisy bus is ON, Cryocooler is OFF with the Heater ON, Scan mirror rotating is OFF and earth shield is CLOSED.
  2. AMSR-E is in Normal mode, rotating at 40 RPM.
  3. AMSU A1 is in Full Scan mode.
  4. AMSU A2 is in Full Scan mode.
  5. CERES Aft is in Crosstrack mode and the covers are CLOSED.
  6. CERES Fore is in Crosstrack mode and the covers are CLOSED.
  7. HSB is in Scan Normal mode and the antenna is in the Target position.
  8. MODIS is outgassing in Standby mode, Nadir Door is Unlatched and CLOSED, Solar Diffuser Door is Unlatched and CLOSED, and the Space View Door is Unlatched and OPEN @ 250 steps (outgassing position).

Activities Deferred:

  1. None


  1. The AMSR-E Signal Processor for the Sensor Unit (SPS) switched off shortly after being commanded to Normal (Science) mode. SPS recycled ON/OFF several times over the weekend. The investigation is currently on-going.

SpaceRef staff editor.